OpenStudy (anonymous):

Show the correct graph of -x-2y=6

OpenStudy (a1234):

You want to get that into the form of y = mx + b. -x - 2y = 6 -2y = x + 6 y = -(x + 6)/2 y = -0.5x + 3

OpenStudy (a1234):

Then you graph it using the y-intercept and slope.

OpenStudy (skullpatrol):

You @a1234 have the wrong sign on 3.